How many grams of K2CrO4 will I get if I react 75g FeCr2O4?

FeCr2O4 + K2CO3 + O2 --> K2CrO4 + Fe2O3 + CO2



1
Expert's answer
2020-12-09T06:36:34-0500

The balanced equation is

4FeCr2O4 + 8K2CO3 + 7O2 --> 8K2CrO4 + 2Fe2O3 + 8CO2


"m(K_2CrO_4)=75g(FeCr_2O_4)\times\frac{1mol(FeCr_2O_4)}{223.83g(FeCr_2O_4)}\times\frac{8mol(K_2CrO_4)}{4mol(FeCr_2O_4)}\times\frac{194.19g(K_2CrO_4)}{1mol(K_2CrO_4)}=130g"


Answer: 130 g